OpenStudy (anonymous):

will give fan and medal solve for x -3x - 8 = 10

pooja195 (pooja195):

Add 8 to both sides 10+8=18 Divide both sides by -3 18/-3=?

OpenStudy (anonymous):

18 \-3x

OpenStudy (anonymous):

x = -6

pooja195 (pooja195):

@bwateam It would have been better if you let the asker find the answer

OpenStudy (anonymous):

im just trying to help free space

OpenStudy (katnissstyles000):

-6

pooja195 (pooja195):

Direct answers are NOT allowed explain your answer.

OpenStudy (katnissstyles000):

ok. Let's solve your equation step-by-step. −3x−8=10 Step 1: Add 8 to both sides. −3x−8+8=10+8 −3x=18 Step 2: Divide both sides by -3. −3x −3 = 18 −3 x=−6 Answer: x=−6

OpenStudy (anonymous):

thanks so much

OpenStudy (katnissstyles000):

welcome

OpenStudy (anonymous):

can you help with more ?

OpenStudy (katnissstyles000):

sure. what do u need help on?

OpenStudy (anonymous):

Evaluate. 

OpenStudy (anonymous):

if you cant help with that one i have more lol @KatnissStyles000

OpenStudy (katnissstyles000):

1 whole? is that what u mean?

OpenStudy (anonymous):

1.6666666667

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so whats thecanswer? @bwateam

OpenStudy (anonymous):

thats the desivmal it should be 1-2\3

OpenStudy (katnissstyles000):

i agree with @bwateam

OpenStudy (anonymous):

@katnissstyles000 thx

OpenStudy (anonymous):

that answer isnt on there

OpenStudy (katnissstyles000):

welcome ;-)

OpenStudy (anonymous):

whats the awnsers

OpenStudy (anonymous):

A) 1 B) 28 C) 4 D) 8

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

ikr lol

OpenStudy (anonymous):

that makes logic

OpenStudy (anonymous):

so do u know the answer or should i just skip this one ?

OpenStudy (anonymous):

i got 2.0001 that time

OpenStudy (anonymous):

so do you think it could be 1

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks can u help with more

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

36) Solve: 3(2x - 5) = 6x - 15 A) 0 B) 1 Eliminate C) 20 D) infinitely many solutions

OpenStudy (anonymous):

6x-15=6x-15

OpenStudy (anonymous):

ims bud

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if u need any help kik me @1stkill

OpenStudy (anonymous):

32) Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C) 1 2 x + 8 = 10 D) 1 2 (2x - 6) = -6

OpenStudy (anonymous):

b

OpenStudy (anonymous):

byr gtg

OpenStudy (anonymous):

okay